OpenStudy (anonymous):

which of the following defines a function f for which f(-x)=-f(x) log(x) sin(x) e^x x^2 cos(x)

OpenStudy (anonymous):

sin(x)

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

f(-x)=-f(x) is the definition of odd function

OpenStudy (anonymous):

so all the other ones are even functions?

OpenStudy (anonymous):

no!... just when f(-x)=-f(x).. for example sin(-2)=-sin(2)=0,035 but e^-2 = 0,13 and -e^2=7,4

OpenStudy (anonymous):

sorry... i describe odd functions but for even functions the answer is no! too

OpenStudy (anonymous):

got it thanks!!

OpenStudy (anonymous):

f(x) = f(-x) for even functions

OpenStudy (anonymous):

in that case... cos(x) and x^2 are even functions